{F} The equation for a displacement 𝑠(π‘š), at a time 𝑑(𝑠) by an object starting at a displacement of 𝑠0 (π‘š), with an initial velocity 𝑒(π‘šπ‘  βˆ’1 ) and uniform acceleration π‘Ž(π‘šπ‘  βˆ’2 ) is: 𝑠 = 𝑠0 + 𝑒𝑑 + 1 2 π‘Žπ‘‘ 2 A projectile is launched from a cliff with 𝑠0 = 30 π‘š, 𝑒 = 55 π‘šπ‘  βˆ’1 and π‘Ž = βˆ’10 π‘šπ‘  βˆ’2 . The tasks are to: a) Plot a graph of distance (𝑠) vs time (𝑑) for the first 10s of motion. b) Determine the gradient of the graph at 𝑑 = 2𝑠 and 𝑑 = 6𝑠. c) Differentiate the equation to find the functions for: i) Velocity (𝑣 = 𝑑𝑠 𝑑𝑑) ii) Acceleration (π‘Ž = 𝑑𝑣 𝑑𝑑 = 𝑑 2 𝑠 𝑑𝑑2 ) d) Use your results from part c to calculate the velocity at 𝑑 = 2𝑠 and 𝑑 = 6𝑠. e) Compare your results for part b) and part d). f) Find the turning point of the equation for the displacement 𝑠 and using the second derivative verify whether it is a maximum, minimum or point of inflection. g) Compare your results from f) with the graph you produced in a).


1
Expert's answer
2022-02-15T16:59:31-0500
"s(t)=30+55t-\dfrac{10t^2}{2}, m"

a)



"s(t)=30+55t-5t^2, 0\le t\le10"


b)



"grad \ s=\dfrac{s_2-s_1}{t_2-t_1}"

"t=2 , s(2)=120"


"t=0 , s=50"



"grad \ s|_{t=2}=\dfrac{120-50}{2-0}=35(m/s)"

"t=6 , s(6)=180"


"t=0 , s=210"

"grad \ s|_{t=6}=\dfrac{180-210}{6-0}=-5(m/s)"


c)

i)



"v(t)=\dfrac{ds}{dt}=55-10t, m/s"

ii)



"a(t)=\dfrac{dv}{dt}=\dfrac{d^2s}{dt^2}=-10\ m/s^2"

d)



"v(2)=55-10(2)=35(m\/s)""v(6)=55-10(6)=-5(m\/s)"

e) The results are the same.


f) The turning point of the equation for the displacement


"\dfrac{ds}{dt}=0=>55-10t=0""t=5.5 s""s(5.5)=30+55(5.5)-5(5.5)^2=181.25(m)"

Turning point isΒ "(5.5, 181.25)"



"\dfrac{d^2s}{dt^2}=-10<0"

PointΒ "(5.5, 181.25)"Β is a maximum.


g) The results are the same